# ChainReaction! Structured Approach with Causal Chains as Intermediate Representations for Improved & Explainable Causal Video Question Answering

### v1

@@ 4 Approach
Thus far, we saw that our approach factorizes VideoQA problem into subproblems: causal chain extraction and causal chain-driven answerer. Now, we explain these individual modules and their training/inference in the following.
Causal Chain Extractor (CCE)
@@
The objective of CCE module is to explicitly capture the cause-and-effect relationships or the reasoning steps from videos, conditioned on the questions, and express them in a detailed yet concise form. Generating causal chains from videos conditioned on questions constitutes a complex reasoning task. Towards that end, we propose to leverage the representational capacity of foundation models. However, generating causal chains is non-trivial task that extends beyond the native capabilities of existing vision-language foundation models. Although foundation models exhibit strong performance across a wide range of vision-language benchmarks, they demonstrate limited proficiency in causal chain generation, primarily due to the absence of structured causal supervision in their pretraining corpora. To address this limitation, a novel dataset comprising video–question–causal chain triplets is introduced, as described in Section 3. This dataset enables the supervised fine-tuning (SFT) of pretrained vision-language foundation models to align their generative capacity with the demands of structured causal inference. Through this process, the model is explicitly trained to capture event-level dependencies and produce human-interpretable causal chains grounded in both visual and linguistic context. Formally, given a video $\mathcal{V}$ and a corresponding question $\mathcal{Q}$, the model learns a mapping: $f_{CCE}:(\mathcal{V},\mathcal{Q})\rightarrow\mathcal{C}$; where $\mathcal{C}$ denotes the generated causal chain and $f_{CCE}$ represents the fine-tuned vision-language foundation model.
Figure 3: Training of both modules in stages.
Causal Chain-Driven Answerer (CCDA)
@@
The objective of CCDA module is to select the correct answer from candidate answers, based on the questions and the causal chains. Existing large language models have shown strong performance on a variety of language-based tasks. We propose leveraging language models to implement the Causal Chain-Driven Answerer (CCDA). However, processing causal chains using language models is non-trivial and mostly not covered in their training suites, largely due to lack of explicit causal chains or step-by-step annotations of causal reasoning. Towards that end, we resort to supervised finetuning of CCDA. Specifically, CCDA receives the extracted causal chain, question, and candidate answer options (typically four to five in existing datasets) and is prompted to select the correct answer. Notably, the Causal Chain Extractor (CCE) remains frozen at this stage. Formally, given a question $\mathcal{Q}$, the corresponding causal chain $\mathcal{C}$, and answer options $\mathcal{O}$ the model learns a mapping: $f_{CCDA}:(\mathcal{Q},\mathcal{C},\mathcal{O})\rightarrow A$; where $\mathcal{A}$ denotes the selected answer option and $f_{CCDA}$ represents a finetuned language model.
Inference
@@
Causality-aware Captioning Metric
To evaluate the quality of generated causal chains, we also assess how respectful the chains are in terms of modeling causality. Since regular captioning metrics do not capture this aspect, we introduce a novel Causal Coherence Score—CauCo score, as an indicator of whether causal chains are causally coherent, i.e., events in the chains are logically connected by cause and effect. To achieve this, following a similar approach of LLM as verifiers, we SFT a LLM to identify if a chain is causally coherent. The final expected output of this evaluator is expected to be ”True” or ”False”. For SFT, we construct a set consisting of positive and negative samples. The positive samples are correct causal chains. To generate negative samples, we perturb the correct chains with a possibility of six distinct strategies, namely: actor swapping, event negation, event removal, event order reversal, semantic modification, and chain shuffling.
@@ 5 Experiments
Implementation details. We use PyTorch (Paszke et al. 2019) to implement models. Noting the strong performance of VILA 1.5 and LLaMA on vision and language tasks, we adopt them as representative models in implementing our CCE and CCDA, respectively. Note that our approach is not designed for or limited to any models; practitioners may use models of their choice such as (Zhang, Li, and Bing 2023; Li et al. 2024; Lin et al. 2024; Chen et al. 2024). Our CCE is based on VILA-3B (Lin et al. 2024); and our CCDA is LLaMA-3.1-8B (Grattafiori et al. 2024) version. Further implementation details are provided in Appendix. We will release the codebase along with causal chains.
Datasets. Unless specified otherwise, we conduct experiments on the datasets as discussed in Section 3.
@@
Performance metric. Following prior work (Xiao et al. 2021; Li, Niu, and Zhang 2022; Parmar et al. 2024), we use Accuracy as the performance metric. For causal chain quality, we use our CauCo score and captioning metrics: BLEU (Papineni et al. 2002), Meteor (Banerjee and Lavie 2005), ROUGE (Lin 2004), SPICE (Anderson et al. 2016) scores.
Experimental Upper bound on Performance
NextQA CVQA CausalChaos! Avg
99.70 99.85 98.65 99.40
@@
Figure 4: Sensitivity to causal chains. (1) plot of performance drop vs links removed. (2) Qualitative example. OCC: original causal chains, MCC: masked chains, SA: selected answer. (3) Qualitative examples of SA changes based on hypothetical chains (HCC). Please zoom-in.
@@
Probing sensitivity to causal chains.
We investigated how the CCDA depended on causal chains; and how sensitive it is to the quality of causal chains. For this, we incrementally mask-out links of the causal chains. Quantitative and qualitative results in Figure 4(1,2) show how the performance degrades as masking is increased. We also tested how hypothetical causal chains influence CCDA’s answer selection in an intuitive way (Figure 4(3)).
Performance Comparison with SOTA
@@
Following prior work (Xiao et al. 2021; Li, Niu, and Zhang 2022; Parmar et al. 2024), we compare against a wide range of models: 1) traditional approaches {BlindQA, EVQA, CoMem, HME, HCRN, HGA}; 2) causal approaches {MCR+HCRN, VCSR}; and 3) SOTA VLMs/Multimodal foundation approaches {MIST, BLIP-2, Video-LLaMA, VideoChat2, GPT-4o, VILA, DeepSeekVL2, Qwen-VL2.5} which excel on vision-language tasks.
Model NXQA CVQA CChQA Avg WtAvg
@@
BlindQA (Antol et al. 2015) 28.38 59.46 13.07 33.64 44.87
EVQA (Antol et al. 2015) 42.31 60.95 13.48 38.91 50.62
CoMem (Gao et al. 2018) 46.15 62.79 13.88 40.94 53.05
HME (Fan et al. 2019) 46.52 61.45 14.02 40.66 52.43
HCRN (Le et al. 2020) 47.00 61.61 17.00 41.86 52.93
HGA (Jiang and Han 2020) 47.00 63.51 15.36 42.00 53.88
@@
VCSR (Wei et al. 2023) 53.00 65.41 - - -
MCR+HCRN (Zang et al. 2023) 49.20 66.00 - - -
@@
BLIP-2 (Li et al. 2023) 45.00 62.00 23.32 43.44 53.00
VideoLLaMA (Zhang, Li, and Bing 2023) 42.00 31.00 11.73 28.24 33.31
VideoChat2 (Li et al. 2024) 60.00 46.00 15.36 40.45 48.50
GPT4o (OpenAI 2024) 70.00 52.00 48.17 56.72 58.00
VILA 1.5-3B (Lin et al. 2024) 60.23 72.11 62.80 65.05 67.20
DeepSeek-VL2 (Wu et al. 2024) 51.55 57.08 17.12 41.92 51.95
QwenVL 2.5-7B (Bai et al. 2025) 70.75 76.22 31.54 59.50 70.73
@@
Table 2: Performance evaluation with SOTA methods. nxqa: NextQA; cvqa: CausalVidQA; cchqa: CausalChaos! datasets. WtAvg: Weighted average.
@@ Results.
Table 2 shows that traditional VideoQA models lag behind VLMs. Notably, MIST—though not a large model—performed strongly, likely due to its long-term video understanding design and use of CLIP features with advanced spatiotemporal attention. Moreover, our method is simpler yet more effective, significantly outperforming prior causal models like MCR+HCRN (best performing model in (Zang et al. 2023))—closest to our work. Additionally, our model provides interpretable, symbolic causal chains—distinct from latent-variable or counterfactual approaches. Next-generation VLMs like (Lin et al. 2024), trained on diverse multimodal tasks, outperformed even the closed-source GPT-4o on causal VideoQA. Building on (Lin et al. 2024), our framework outperforms all existing models, likely due to: 1) improved visual grounding via causal chain extraction; and 2) explicit reasoning in the CCDA. In contrast, VLMs perform grounding and reasoning jointly, which may cause them to overlook visual evidence and rely on biases like plausibility or correlation. See Figure 3 for qualitative results. Our model takes a structured approach, guiding the causal chain extractor to capture cause-and-effect relationships in the video. We view this as a dense prediction task, where fine-grained supervision enhances video understanding. In contrast, typical VideoQA models use a single-label objective, which can lead to shallow learning—guessing answers without true video understanding or causal reasoning. Our approach avoids this by isolating video comprehension from bias, allowing the CCDA to focus solely on selecting the correct answer from the extracted causal chains (reasoning steps). Causal chains provide concise, focused context. While the CCDA made fewer errors, the Causal Chain Extractor sometimes misinterpreted object roles or actions despite correct identification, leading to flawed chains and reduced CCDA accuracy (see Figure 4). A good strategy for future research might be pursue improvements in: 1) role or relationship modeling; and 2) situation understanding. We further compare the quality of generated causal chains from our model with a representative SOTA VLM-QwenVL2.5 in Table 3—results averaged across all datasets. We found that our model performs significantly better than VLM counterpart. This indicates that causality aspect maybe overlooked in the current multimodal understanding literature (highlighted in (Ma 2025)), a gap that our work fills. We believe models like ours also have the potential to serve as reasoning engines, which should be explored by future work.
@@
Human Studies
Explainable systems benefit two key user groups: researchers/system designers and consumers, each with distinct objectives. We outline their goals and systematically design human subject studies to evaluate how our explainable system compares to black-box models across various criteria. Most SOTA models/VLMs operate as black boxes. We select VILA (Lin et al. 2024) as a representative black-box model due to its strong performance in previous experiments and lack of explicit model explanations. To minimize bias, all the participants of this study were not a part of the project. Following previous study (Parmar et al. 2024), we perform all user studies with six participants—one more than the previous study.
Study I: Explainability
@@
Study III: Human Preference
@@
Study IV: Utility in System Debugging
@@
Researchers and system designers, responsible for developing AI systems, aim to evaluate and improve model performance through systematic debugging. Identifying limitations is a key first step, where intermediate outputs—alongside final predictions—offer valuable insights. To test this hypothesis, we designed an experiment targeting failure cases of both our explainable and black-box models. Although it assumes knowledge of incorrect predictions, this reflects a realistic scenario, as system designers typically work with labeled data. Researchers are then asked to identify the source of the model’s failure: 1) visual perception, 2) language and reasoning, or 3) Cannot Tell. The third option indicates uncertainty, while choosing the first two suggests a clearer understanding of the system and supports more confident diagnosis. Six researchers with expertise in computer vision, VLMs, and LLMs each analyze 20 failure cases from both systems, identifying fault locations. For our model, they view the video, question, causal chain explanations, and final predictions; for the black-box model, only the video, question, and final predictions are provided. Participants are instructed to use all available information in their assessments. Table 5 shows that the ”Cannot Tell” option was chosen far less often for our model than in case of BlackBox model. Researchers attributed failures to the chain extractor in $\sim$48% of cases and to the chain-driven answerer in $\sim$38%, suggesting that causal chains improve system debuggability.
Out-Of-Domain Chain Generation
@@
To test the generalizability of CCE, we train it on a cartoon-based dataset (CausalChaos!) and evaluate it on an out-of-distribution real-world dataset (NextQA), making the task intentionally challenging. Quality of generated chains is evaluated in terms of bleu-1–4 (b), meteor (m), rouge (r), spice (s), our causal coherence score ccs, with groundtruth causal chains (Section 4). A zero-shot vila 1.5 serves as the baseline; our extractor also uses vila 1.5 for a fair comparison. Table 6 shows that our model significantly outperforms the baseline in cross-dataset causal chain generation—suggesting that the reasoning patterns that our causal chain extractor learns are robust and transferable, highlighting its potential as a reusable causal reasoning engine. Qualitative results in Fig 6.

### v2

@@ 4.2 Model
Thus far, we saw that our approach factorizes VideoQA problem into modules: causal chain extraction and causal chain-driven answerer. Now, we explain these individual modules and their training/inference in the following.
4.2.1 Causal Chain Extractor (CCE)
@@
The objective of CCE module is to explicitly capture the cause-and-effect relationships or the reasoning steps from videos, conditioned on the questions, and express them in a detailed yet concise form. The CCE extracts the causal chain after the entire video has been observed, when the full causal process and its outcome are known. At this stage, the causal structure is realized as a single factual sequence of events leading to the answer. The extractor therefore identifies this linear causal path directly, without reasoning over counterfactual or hypothetical alternatives. This formulation treats causal chain extraction as a descriptive recognition task, emphasizing interpretability, fidelity, and robustness within the VideoQA paradigm. Generating causal chains from videos conditioned on questions is a complex reasoning task. We leverage the representational capacity of foundation models. However, generating causal chains is non-trivial and beyond existing vision-language foundation models. Although foundation models exhibit strong performance across vision-language benchmarks, they lack structured causal supervision, limiting causal chain generation. To address this limitation, we introduce a novel dataset of video–question–causal chain triplets (Section 3). This dataset enables supervised fine-tuning (SFT) of pretrained vision-language models for structured causal inference. This trains the model to capture event-level dependencies and generate human-interpretable causal chains grounded in visual and linguistic context. Formally, given a video $\mathcal{V}$ and a corresponding question $\mathcal{Q}$, the model learns a mapping: $f_{CCE}:(\mathcal{V},\mathcal{Q})\rightarrow\mathcal{C}$; where $\mathcal{C}$ denotes the generated causal chain and $f_{CCE}$ represents the fine-tuned vision-language foundation model. Further model implementation details provided in Appendix.
Figure 3: Stage-wise training of our model.
4.2.2 Causal Chain-Driven Answerer (CCDA)
@@
The objective of CCDA module is to select the correct answer from candidate answers, based on the questions and the causal chains. Existing large language models have shown strong performance on a variety of language-based tasks. We propose leveraging language models to implement the CCDA. However, processing causal chains using language models is non-trivial and mostly not covered in their training suites, largely due to lack of explicit causal chains or step-by-step annotations of causal reasoning. Towards that end, we resort to supervised finetuning of CCDA. Specifically, CCDA receives the extracted causal chain, question, and candidate answer options (typically four to five in existing datasets) and is prompted to select the correct answer. Notably, the Causal Chain Extractor (CCE) remains frozen at this stage. Formally, given a question $\mathcal{Q}$, the corresponding causal chain $\mathcal{C}$, and answer options $\mathcal{O}$ the model learns a function: $f_{CCDA}:(\mathcal{Q},\mathcal{C},\mathcal{O})\rightarrow\mathcal{A}$; where $\mathcal{A}$ denotes the selected answer option and $f_{CCDA}$ represents a finetuned language model.
4.2.3 Clean Separation with Stage-wise Training
Problem with end-to-end (E2E) training. E2E VQA models optimize all components—from feature extraction to answer prediction—using final answer loss. This causes gradient leakage, letting error signals flow through the entire pipeline. As a result, models may learn “shortcuts” that boost final answer accuracy but harm causal reasoning.
Our solution—stagewise training. Our two-stage, stagewise training prevents this by introducing a clear boundary:
  • CCE Training: The CCE is trained independently to match the ground-truth causal chains ($C$), using a loss defined on $C$ that enforces accurate causal grounding.
  • CCDA Training: The CCDA is trained on the CCE’s output, with the CCE’s weights typically frozen or its predictions treated as fixed during this stage.
By freezing the CCE, no gradients can pass backward from the CCDA’s answering loss to the CCE’s internal weights. This ensures that the CCE’s learned causal reasoning logic remains pure and untainted by the pressure to maximize the answer score, thus preserving the integrity, robustness, and causal semantics of the intermediate representation.
4.2.4 Inference
@@
4.3 CauCo: A Causal Coherence Metric
We introduce the CauCo score to evaluate how well generated causal chains model causality—an aspect regular captioning metrics miss. CauCo quantifies causal consistency and provides causal guarantees in open-world settings by measuring whether events are logically linked by cause and effect. Following the LLM-as-verifier approach, we SFT an LLM to judge whether a chain is causally coherent. The evaluator outputs “True” or “False.” For SFT, we construct positive and negative samples: positive ones are correct causal chains, and negatives are created by perturbing them using six strategies—actor swapping, event negation, event removal, event order reversal, semantic modification, and chain shuffling.
@@ 5 Experiments
Implementation details. We use PyTorch [paszke2019pytorch] to implement models. Noting the strong performance of VILA 1.5 and LLaMA on vision and language tasks, we adopt them as representative models in implementing our CCE and CCDA, respectively. Note that our approach is not designed for or limited to any models; practitioners may use models of their choice such as [videollama, videochat2, vila, internvl25]. Our CCE is based on VILA-3B [vila]; CCDA is LLaMA-3.1-8B [grattafiori2024llama] version. Further implementation details are provided in Appendix. Codebase along with causal chains will be released.
Datasets. We conduct experiments on the datasets as discussed in Section 3.
@@
Performance metric. Following prior work [nextqa, causalvidqa, causalchaos], we use Accuracy as the performance metric. For causal chain quality, we use our CauCo score and captioning metrics: BLEU [bleu], Meteor [meteor], ROUGE [rouge], SPICE [spice].
5.1 Experimental Upper bound on Performance
Dataset NextQA CVQA CausalChaos! Overall
CCDA Accuracy 99.70 99.85 98.65 99.40
@@ 5.1 Experimental Upper bound on Performance
Figure 4: Causal chain ablation study. (1) Study-I: QA accuracy drops by 73% when chains are perturbed. (2) Study-II: drop in QA is correlated to amount of perturbation. (3,4) Qualitative example. OCC: original causal chains, MCC: masked chains, SA: selected answer, PCC: perturbed chains. SA changes intuitively as chains are perturbed. Please zoom-in.
5.2 Ablation Studies on the Role of Causal Chains
To test whether CCDA’s improvements arise from genuine causal reasoning rather than surface-level context enrichment, we conducted ablation studies that systematically degrade the quality of causal chains. Study-I: We semantically perturbed the causal chains such that contextual information was preserved but causal relations were disrupted. We observed a sharp decline in QA accuracy once causal information was perturbed, despite the context remaining intact (Figure 4(1)). Study-II: We progressively masked links and measured QA performance; accuracy degraded monotonically with increased perturbation (Figure 4(2)). Causal-chain quality strongly correlated with CCDA accuracy (r=0.97). These results empirically validate that CCDA’s reasoning depends on causal-chain integrity, with improvements driven by causal reasoning rather than spurious correlations. Qualitative examples (Figure 4(3,4)) show how chain perturbations intuitively alter CCDA’s answers.
5.3 Performance Comparison with SOTA
@@
Following prior work [nextqa, causalvidqa, causalchaos], we compare against a wide range of models: 1) traditional approaches [evqa, comem, hme, hcrn, hga]; 2) causal approaches [wei2023visual, zang2023discovering]; and 3) SOTA VLMs/Multimodal foundation approaches [mist, blip2, videollama, videochat2, gpt4o, vila, deepseekvl2, Qwen2.5-VL], which excel on vision-language tasks.
Model NeXT-QA [nextqa] CausalVidQA [causalvidqa] CausalChaos!QA [causalchaos] Avg WtAvg
@@
BlindQA [evqa] 28.38 59.46 13.07 33.64 44.87
EVQA [evqa] 42.31 60.95 13.48 38.91 50.62
CoMem [comem] 46.15 62.79 13.88 40.94 53.05
HME [hme] 46.52 61.45 14.02 40.66 52.43
HCRN [hcrn] 47.00 61.61 17.00 41.86 52.93
HGA [hga] 47.00 63.51 15.36 42.00 53.88
@@
VCSR [wei2023visual] 53.00 65.41 - - -
MCR+HCRN [zang2023discovering] 49.20 66.00 - - -
@@
BLIP-2 [blip2] 45.00 62.00 23.32 43.44 53.00
VideoLLaMA [videollama] 42.00 31.00 11.73 28.24 33.31
VideoChat2 [videochat2] 60.00 46.00 15.36 40.45 48.50
GPT4o [gpt4o] 70.00 52.00 48.17 56.72 58.00
VILA 1.5-3B [vila] 60.23 72.11 62.80 65.05 67.20
DeepSeek-VL2 [deepseekvl2] 51.55 57.08 17.12 41.92 51.95
QwenVL 2.5-7B [Qwen2.5-VL] 70.75 76.22 31.54 59.50 70.73
@@
Table 2: Performance evaluation with SOTA methods. WtAvg: Weighted average.
@@ Results.
Table 2 reveals that traditional VideoQA models lag behind VLMs. MIST, a smaller model, performs well—likely from its long-term video understanding and use of CLIP features with spatiotemporal attention. Our method further outperforms prior causal models such as MCR+HCRN [zang2023discovering], the closest to our work, while remaining simpler and producing interpretable causal chains, unlike latent-variable or counterfactual methods. Next-generation VLMs such as VILA [vila] outperform GPT-4o on causal VideoQA. Building on VILA, our framework surpasses prior models through improved visual grounding via causal chain extraction and explicit CCDA reasoning. In contrast, VLMs perform grounding and reasoning jointly, often overlooking visual evidence and relying on plausibility biases. Our structured model guides the extractor to capture cause–effect relations in videos—a dense prediction task enhanced by fine-grained supervision. Typical VideoQA models use a single-label objective, leading to shallow learning, whereas our approach isolates comprehension from bias, allowing the CCDA to focus on selecting answers from extracted causal chains. The CCDA made fewer errors, though the extractor occasionally misinterpreted object roles or actions, producing flawed chains and reducing accuracy (Figure 5). Based on our analysis, future work should focus on 1) role or relationship modeling and 2) situation understanding. We also compare our model’s causal chains with those from QwenVL2.5, a representative SOTA VLM (Table 3, averaged across datasets). Our model performs significantly better, suggesting that causality is often overlooked in multimodal understanding [ma2025causal]—a gap our work addresses. We believe models like ours also have the potential to serve as reasoning engines, which should be explored by future work.
@@
5.4 Human Studies
Explainable systems benefit two user groups: researchers/ system designers & consumers. We design four studies to evaluate our explainable system against black-box models across multiple criteria. Since most SOTA VLMs are black boxes, we use VILA [vila] as a representative model for its strong performance & lack of explicit explanations. To ensure reliability, we conducted user studies with six participants, exceeding the sample size used in prior work [causalchaos]. To minimize bias: 1) no study participants were involved in the project; 2) Sample order randomization & blinding (independent, anonymized tests) was used for all participants. Further details provided in Appendix.
5.4.1 Study I: Explainability
@@ 5.4.2 Study II: Trustworthiness
5.4.3 Study III: Human Preference
@@
5.4.4 Study IV: Utility in System Debugging
@@
Researchers and system designers seek to improve models through systematic debugging. Identifying limitations is the first step; intermediate outputs such as causal chains alongside final predictions provide useful insights. We tested this by examining failure cases of our explainable and black-box models. While this assumes known errors, it reflects real settings where designers work with labeled data. Researchers are then asked to identify the source of the model’s failure: 1) visual perception, 2) language and reasoning, or 3) Cannot Tell. The third option indicates uncertainty, while choosing the first two suggests a clearer understanding of the system and supports more confident diagnosis. Six researchers with expertise in computer vision, VLMs, and LLMs each analyze 20 failure cases from both systems, identifying fault locations. For our model, they view the video, question, causal chain explanations, and final predictions; for the black-box model, only the video, question, and final predictions are provided. Participants are instructed to use all available information in their assessments. Table 5 shows that the ”Cannot Tell” option was chosen far less often for our model than in case of BlackBox model. Researchers attributed failures to the chain extractor in $\sim$48% of cases and to the chain-driven answerer in $\sim$38%, suggesting that causal chains improve system debuggability.
5.5 Out-Of-Domain Chain Generation
@@
To test the generalizability of CCE, we train it on a cartoon-based dataset (CausalChaos!) and evaluate it on an out-of-distribution real-world dataset (NextQA), making the task intentionally challenging. Quality of generated chains is evaluated in terms of bleu-1–4 (b), meteor (m), rouge (r), spice (s), our causal coherence score ccs, w.r.t. groundtruth causal chains (Section 3). A zero-shot vila 1.5 serves as the baseline; our extractor also uses vila 1.5 for a fair comparison. Table 6 shows that our model significantly outperforms the baseline in cross-dataset causal chain generation—suggesting that the reasoning patterns that our causal chain extractor learns are robust and transferable, highlighting its potential as a reusable causal reasoning engine. Qualitative results in Figure 6.